OpenStudy (anonymous):

7x = 20 + 4x It says my answer options are : (A) 5 and 2/3 , (B) 6 and 1/3 , (C) 6 and 2/3 , and (D) 28?

OpenStudy (anonymous):

3x=20 3x-20=0 7x=4(x+5) x=20/3

OpenStudy (anonymous):

why is 6= 2/3 the same as 20/3?

OpenStudy (anonymous):

Even I have no clue

OpenStudy (anonymous):

the solution to that problem was not given.

OpenStudy (anonymous):

I mean how is 6= 2/3 =20/3 ?

OpenStudy (anonymous):

7(20/3) = 20 + (4(20/3)) 46.667 = 46.667

OpenStudy (anonymous):

ohhhhhh

OpenStudy (anonymous):

thank u I got it now

OpenStudy (anonymous):

\[x - 2 (5/7) = (2/3) \]

OpenStudy (anonymous):

x−2(5/7)=(2/3)